OpenStudy (anonymous):

I really don't understand. Any inputs? Simplify. Express the product as a radical expression. 11√x^5 times √x

OpenStudy (ash2326):

We have \[11 \sqrt {x^5} \times \sqrt x\] We know that \[\sqrt A \times \sqrt B=\sqrt{AB}\] Can you use this here?

OpenStudy (anonymous):

My answer: 

OpenStudy (ash2326):

First step multiply these \[\sqrt{x^5}\times \sqrt x=???\]

OpenStudy (anonymous):

you're doing it a different way

OpenStudy (ash2326):

How did you approach the problem? @InsanelyChaotic

OpenStudy (anonymous):

Am I wrong? @mark_o.

OpenStudy (anonymous):

\[11\left( x ^{5} \right)^{\frac{ 1 }{ 2 }}\left( x \right)^{\frac{ 1 }{ 2 }}\] \[=11\left( x \right)^{\frac{ 5 }{ 2 }}\left( x \right)^{\frac{ 1 }{ 2 }}\] add the power of x, then

OpenStudy (anonymous):

I'm sorry, I don't understand

OpenStudy (anonymous):

\[=11\left( x \right)^{\frac{ 6 }{ 2 }}\] divide the power to get \[=11\left( x \right)^{3}\] ................did you follow this?

OpenStudy (anonymous):

= 1

OpenStudy (anonymous):

radical expressions can also be written as using the power function so that you can easilyadd or subtract powers \[11\sqrt{x ^{5}}*\sqrt{x}\] \[=11\left( x ^{5} \right)^{\frac{ 1 }{ 2 }}*\left( x \right)^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

Do I add both halfs?

OpenStudy (anonymous):

halves*

OpenStudy (anonymous):

yes you can

OpenStudy (anonymous):

1/2 + 1/2 = 1 whole

OpenStudy (anonymous):

its the power of \[\frac{ 5 }{ 2 }+\frac{ 1 }{ 2 }=\frac{ 6 }{ 2 }=3\]

OpenStudy (anonymous):

\[=11\left( x ^{5} \right)^{\frac{ 1 }{ 2 }}*\left( x \right)^{\frac{ 1 }{ 2 }}\] multiply first the power of 5 and 1/2, to get \[=11\left( x \right)^{\frac{ 5 }{ 2 }}*\left( x \right)^{\frac{ 1 }{ 2 }}\] now add the powers of both x \[=11\left( x \right)^{\frac{ 6 }{ 2 }}\] dividing the power \[=11\left( x \right)^3\] \[=11x ^{3}\]

OpenStudy (anonymous):

the rule on powers is to multiply first then add right?